4)An inflated balloon has a colume of 6.0l sea level (1.0atm) and is allowed to ascend in altitude until the pressure is 0.45 ATM. during ascent the temp of the gas falls from 22 c to -21 c. calculate the volume of the balloon at its final volume?

5)A can of hair spray should never be incinerated. in addition to the contents being flammable, the pressurized is danger itself. if the gas pressure in an aerosol can is 1.5 atm at 25c. what would be the prssure if the can was heated to 450c ( you may assume the gas inside the can obeys the ideal-gas equation.)

6) a gaseous mixture made of 6.00g of 02 and 9.00g of CH4 is placed
1
Expert's answer
2019-11-12T12:02:11-0500

1) Use Combined Law:


"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"

"T_1 (K) = 22+273 = 295 K"

"T_2(K) = -22 +273 = 251 K"



"\frac{1.0\times6.0}{295} = \frac{0.45\times V_2}{251}"

"V_2 = 11.3 l"

2) Use Gay-Lussac's Law:


"\frac{P_1}{T_1}=\frac{P_2}{T_2}"

"T_1(K)= 25+273 = 298 K"

"T_2(K) = 450+273 = 723 K"


"\frac{1.5 }{298} =\frac{P_2}{723}"

"P_2 =3.6 atm"

3) not enough data